A beaker with 165 mL of an acetic acid buffer with a pH of 5.000 is sitting on a benchtop. The total molarity of acid and conjugate base in this buffer is 0.100 M. A student adds 4.00 mL of a 0.360 M HCl solution to the beaker. How much will the pH change? The pKa of acetic acid is 4.740.
Answer #1

According to Henderson-Hesselbalch equation pH pKa+ log (base/acid) let x concentration of acid let y concentration of conjugChanges after addition of HCl: moles conjugate base = 0.01043-0.00144-0.00899 moles acid = 0.006072 + 0.00144=0.00751 total v